OpenStudy (magbak):

I WILL AWARD MEDAL!!!!!!!!!!!!!!! PLEASE HELP!!!!!!!!!!!!!!!!!!!!! Write log8 3 as a logarithm of base 5.

OpenStudy (magbak):

@souvik @whpalmer4 @uri

OpenStudy (souvik):

\[\log_{8}3 \] right?

OpenStudy (magbak):

yes

OpenStudy (magbak):

Are you going to help.

OpenStudy (souvik):

\[\log_{a}b=\log_{c}b \times \log_{a}c \]

OpenStudy (souvik):

so.. \[\log_{8}5=\log_{5}5 \times \log_{8}5 \] =\[\frac{ 1 }{ \log_{5}8 }\]

OpenStudy (magbak):

These are the options log base 3 of 5 over log base 8 of 5 log base 5 of 3 over log base 5 of 8 log base 8 of 5 over log base 3 of 5 log base 5 of 8 over log base 5 of 3

OpenStudy (magbak):

Ok thank you I go the correct answer.